Razor:

Solve for x. 2 1/2x−3/4(2x+5)=3/8

Razor:

@563blackghost pleaseeeeeeeeeeee Your great at this stuff ;-;

Razor:

You're*

563blackghost:

Is that by the power?

Razor:

1 attachment
Razor:

^ Does that make it easier to read?

563blackghost:

much better :3

563blackghost:

First yew need to distribute the \(\bf{-\frac{3}{4}}\) into the paranthesis. \(\large\bf{-\frac{3}{4}(2x+5) \rightarrow (2x \times \color{red}{-\frac{3}{4}})+(5 \times \color{red}{-\frac{3}{4}})}\) What does that simplify to?

Razor:

Simplifying -3/4?

Razor:

0.75 but with a negative sign right?

563blackghost:

ye though its best to keep in fraction form

Razor:

its like 0.75 = 75/100 = 3/4

563blackghost:

mhm just simplify. \(\large\bf{-\frac{3}{4} \times 2x =?}\)

Razor:

er -3x/2

Razor:

( ^ I suck at this.. ^^)

silvernight269:

No u don’t. Ur doing fine

Razor:

lol thanks

563blackghost:

you did it right cx

Razor:

@563blackghost miss ghost ? XD

Razor:

xc

563blackghost:

Now simplify the next part? \(\large\bf{5 \times -\frac{3}{4}}\)

silvernight269:

\[-\frac{ 3 }{ 2}x\] Actually

Razor:

3.75?

Razor:

And yes thanks silver xD

563blackghost:

well \(\large\bf{-\frac{3}{2}x \rightarrow -\frac{3x}{2}}\)

563blackghost:

if you were to simplify the fraction...@silvernight269

silvernight269:

-1.5x

563blackghost:

Well -3.75 but correct. You want to keep this as a fraction (due to your answers being in fraction form). So you have... \(\large\bf{-\frac{3}{2}x -3\frac{3}{4}}\) So you have now.. \(\large\bf{2 \frac{1}{2}x -\frac{3}{2}x-3\frac{3}{4}=\frac{3}{8}}\) Now combine like terms... \(\large\bf{2\frac{1}{2}x-\frac{3}{2}x =?}\)

Razor:

How do I do that.. -im soo bad at this type of mathhhh-

563blackghost:

You can convert the fraction into improper. \(\large\bf{2\frac{1}{2} \rightarrow \frac{5}{2}}\) \(\Large\bf{\frac{5}{2}x-\frac{3}{2}x=?}\)

silvernight269:

Just make it into a fraction. Not improper fraction

563blackghost:

\(\Large\bf{\frac{5}{2}x-\frac{3}{2}x ~~\rightarrow~~\frac{5-3}{2}x=?}\) just incase you still dun understand.

Razor:

xD Yeah I might have missed this class.. so I'll try my best to understand -takes out paper- ONWARD

Razor:

5-3=2 so 2/2x?

563blackghost:

Correct. Though simplify further... \(\large\bf{\frac{2}{2} = 1}\) So... \(\large\bf{\frac{5}{2}x-\frac{3}{2}x=\color{red}{x}}\) So you now have... \(\Large\bf{x-3\frac{3}{4}=\frac{3}{8}}\) Follow?

Razor:

Is my answer 3/8? Or are we still going through

563blackghost:

We still going, we need to simplify down so that x is alone.

Razor:

Oof..

563blackghost:

but we just need about 3 more steps then we are done cx

563blackghost:

Now you see that on the left side the fraction \(\large\bf{-3\frac{3}{4}}\) has a denominator of 4. And the fraction on the right; \(\large\bf{\frac{3}{8}}\), has a denominator of 8. In order to get \(\large\bf{-3\frac{3}{4}}\) to the right we need to have like denominators. So we need to first change \(\large\bf{-3\frac{3}{4}}\) into an improper fraction. So how would you change \(\large\bf{-3\frac{3}{4}}\) into improper?

Razor:

uhhhhhh........

Razor:

divide?

563blackghost:

Changing into improper fraction you would multiply the whole number with the denominator and then add it with the numerator. Example: 

563blackghost:

So how would u do it? Changing \(\Large\bf{-3\frac{3}{4}}\)

Razor:

OH KFC Right! KCF thingy

Razor:

It's.. uh 5/8 ....

Razor:

?

silvernight269:

-15/4

563blackghost:

Correct. Now we need to have it a denominator of 8. So we follow by: \(\Large\bf{-\frac{15}{4} \times \frac{2}{2}=\frac{?}{8}}\)

Razor:

Would I do 15 like -15 or just keep it positive? and would I do 2x2 on the top and then 2x2 x 15?

563blackghost:

keep with negative. \(\Large\bf{\frac{-15 \times 2}{4 \times 2} = \frac{?}{8}}\)

Razor:

-30/8

563blackghost:

Gud. So we now have. \(\Large\bf{x-\frac{30}{8}=\frac{3}{8}}\) Now we bring \(\large\bf{-\frac{30}{8}}\) to the right side by `ADDING` (since it is being subtracted with x). \(\Large\bf{x=\frac{3}{8} + \color{red}{\frac{30}{8}}}\) Simplify.

Razor:

33/16

Razor:

And then I simplify that?

563blackghost:

keep the denominator the same

Razor:

33/8*

563blackghost:

when adding fraction you only add the numerator, not the denominator

563blackghost:

Ye right, now you just need to convert to mixed fraction. Can you do that?

Razor:

bleh this is so much work but OKI XD

563blackghost:

ish da last step cx

Razor:

sooo 3 3/8?

563blackghost:

mm not quite.

Razor:

4 1/8?

563blackghost:

Correct!

563blackghost:

So thats your answer cx

Razor:

Finally XD I wish it was day time so I could've finished this faster and not waste your time

563blackghost:

its okie cx